Find T1 for the given isomorphism T. T: P2 → p2 with Tax2 + bx + c) = cx2 – bx + a T ax? + bx + c) =
Answer #1

G OTO (axt bu+C) = (x² - 6x +a T(i) = 0.1 +0.4+10? T(x) = 0.1-1.2 + O. nr T(4²) = 1 +0.1 +0.2² ro o 17-A A/ T = 0 | o lan loo